OpenStudy (anonymous):

Point D(5, -1) is the midpoint of segment CE. If point C has coordinates (3, 2), thenwhat are the coordinates of point E

OpenStudy (anonymous):

@sarah786

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

Im really sorry . Im not good in maths :/

OpenStudy (anonymous):

xE = xD + (xD - xC) = ... + (5 - 3) = ..... yE = yD + (yD - yC) = ... + (-1 - 2) = .....